### v1

@@ Abstract
The $D_{\alpha}$ spectral radius of $G$ is the spectral radius of $D_{\alpha}(G)$, denoted by $\mu_{\alpha}(G)$. In this paper, we establish a lower bound on the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius of non-transmission regular graphs, and we also characterize the extremal graphs attaining the bound.
Key Words:  Generalized distance matrix, Spectral radius, Non-transmission regular graph, Lower bound. AMS Subject Classification (2020):  05C50,15A18
@@ 1 Introduction
Throughout this paper all graphs considered are simple and connected. Let $G=(V(G),E(G))$ be a graph with vertex set $,V(G)=\{v_{1},v_{2},$$\ldots,v_{n}\}$ and edge set $E(G)$. The number of vertices $|V(G)|=n$ is the order of $G$ and the number of edges $|E(G)|=m$ is the size of $G$. For any vertex $v_{i}\in V(G)$, $N(v_{i})=\{v_{j}:(v_{i},v_{j})\in E(G)\}$ is called the set of neighbors of $v_{i}$. The degree $d(v_{i})$ of a vertex $v_{i}$ in $G$ is the number of edges of $G$ incident with $v_{i}$, i.e. , $d(v_{i})=|N(v_{i})|$. Let $K_{n}$ denote the complete graph of order $n$. For other concepts, please refer to [4].
@@
The distance between two vertices $v_{i}$ and $v_{j}$, denoted by $d(v_{i},v_{j})$ or simply by $d_{ij}$, is the length of a shortest path from $v_{i}$ to $v_{j}$ in $G$. The diameter of $G$, denoted by $diam(G)$, is the maximum distance between any two vertices of $G$. For any vertex $v_{i}\in V(G)$, the transmission of $v_{i}$ in $G$, denoted by $Tr_{i}$, is the sum of distances from $v_{i}$ to all other vertices of $G$, that is, $Tr_{i}=\sum\limits_{k=1}^{n}d_{ik}$. For a connected graph $G$, if $Tr_{1}=Tr_{2}=\cdots=Tr_{n}=r$, we say that $G$ is $r$-transmission regular; otherwise, $G$ is non-transmission regular. The second transmission degree of $v_{i}$, denoted by $T_{i}$, which is given by $T_{i}=\sum\limits_{t=1}^{n}d_{it}Tr_{t}$. The sum of all distances between all unordered pairs of vertices in $G$ is called the Wiener index of graph $G$, denoted by $W(G)$, i.e., $W(G)=\frac{1}{2}\sum\limits_{i=1}^{n}Tr_{i}$.
@@
For a connected graph $G$, the distance matrix of $G$ denoted by $D(G)=(d_{ij})_{n\times n}$, where $d_{ij}=d(v_{i},v_{j})$ is the distance from vertex $v_{i}$ to vertex $v_{j}$. Let $Tr(G)=\textrm{diag}(Tr_{1},Tr_{2},\ldots,Tr_{n})$ be the diagonal matrix with vertex transmissions of $G$. With the development of the matrix theory of graphs, Aouchiche and Hansen [1] defined the distance signless Laplacian matrix of $G$ as $D^{Q}(G)=Tr(G)+D(G)$. The eigenvalues of $D(G)$ ($D^{Q}(G)$) are called the distance (distance signless Laplacian) eigenvalues of $G$, which are all real. The largest distance (distance signless Laplacian) eigenvalue of $G$ is called distance (distance signless Laplacian) spectral radius of $G$, denoted by $\mu(G)$ ($\mu_{Q}(G)$).
@@
So far, the distance spectral radius and distance signless Laplacian spectral radius of a connected graph have been investigated extensively. For example, in [18], Nath and Paul determined the unique graph with minimum distance spectral radius among all connected bipartite graphs of order $n$ with a given matching number. In [12], Lin and Lu determined the graphs with minimum and maximum distance signless Laplacian spectral radius among connected graphs with fixed order and clique number. In [14], Lin and Zhou determined the unique trees with second and third maximum distance signless Laplacian spectral radius, respectively, and they also determined the unique connected graph with second maximum distance signless Laplacian spectral radius. In [24], Zhang et al. determined the minimum distance spectral radius of graphs among n-vertex graphs with given connectivity and independence number, and characterized the corresponding extremal graph. In [16], Liu et al. posed a conjecture about the difference between the maximum vertex transmission and the distance spectral radius. In [15], Liu et al. gave a lower bound for the difference between the maximum vertex transmission and the distance spectral radius. In [11], Lan gave a sharp lower bound of $2Tr_{max}-\mu_{Q}(G)$ among all $n$-vertex connected graphs, and characterized the extremal graphs. For more results about distance and distance signless Laplacian spectral radius of graphs, readers may refer to [2, 3, 9, 20, 21, 23] and the references therein.
@@
In [6], Cui et al. proposed to study the convex linear combinations of $Tr(G)$ and $D(G)$, deonted by $D_{\alpha}(G)$ and defined as
@@
We call $D_{\alpha}(G)$ as the generalized distance matrix of $G$. It is easy to see $D_{0}(G)=D(G)$, $D_{1}(G)=Tr(G)$ and $2D_{\frac{1}{2}}(G)=D^{Q}(G)$, thus, the family $D_{\alpha}(G)$ extends both $D(G)$ and $D^{Q}(G)$. Therefore, it is very important to study the generalized distance matrix. The largest eigenvalue of $D_{\alpha}(G)$ is called the $D_{\alpha}$ spectral radius or generalized distance spectral radius of $G$, denoted by $\mu_{\alpha}(G)$.
Recently, the generalized distance spectral radius have been studied in some papers. In [7], Cui et al. characterized the unique graph with minimum generalized distance spectral radius among the connected graphs with fixed chromatic number. In [8], Diaz et al. presented some extremal results on the spectral radius of generalized distance matrix that generalize previous results on the spectral radii of the distance matrix and distance signless Laplacian matrix. In [13], Lin et al. gave several graph transformations on the $D_{\alpha}$-spectral radius. There are also some other interesting results on the $D_{\alpha}$ spectral radius which can be found in [5, 8, 17, 19].
In the theory of graph spectra, there has a classic problem which is establish upper or lower bounds for related spectral radius of given class of graphs. Motivated by [15] and [11], in this paper we decide to study the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius of connected graphs, we establish a lower bound on the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius of non-transmission regular graphs, and we also characterize the extremal graphs attaining the bound.
2 Lower bound on the difference between the maximum vertex transmission degree and the $D_{\alpha}$ spectral radius
In this section, we will establish a lower bound on the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius for graph which is non-transmission regular. First of all, we will give some concepts and lemmas.
@@
(Perron-Frobenius) ([10]) Let $M$ be irreducible and nonnegative square matrix of order $n$. Let $\rho(M)$ be the spectral radius of $M$. Then the following results hold.
@@
Then we give the definition of the equitable quotient matrix and its some properties.
@@
where the blocks $M_{ij}$ are $n_{i}\times n_{j}$ matrices for any $1\leq i,j\leq t$ and $n=n_{1}+n_{2}+\cdots+n_{t}.$ For $1\leq i,j\leq t$, let $b_{ij}$ denote the average row sum of $M_{ij}$, i.e. $b_{ij}$ is the sum of all entries in $M_{ij}$ divided by the number of rows. Then $B(M)=(b_{ij})$(or simply $B$) is called the quotient matrix of $M$. If, in addition, for each pair $i$, $j$, $M_{ij}$ has a constant row sum, then $B$ is called the equitable quotient matrix of $M$.
@@
([22]) Let $B$ be the equitable quotient matrix of $M$. In addition, let $M$ be a nonnegative matrix. Then the spectral radius relation satisfies $\rho(B)=\rho(M)$.
For a connected graph $G$ with order $n$, take $\tau(G)=\frac{Tr_{max}(G)-\mu_{\alpha}(G)}{1-\alpha}$, where $Tr_{max}$ is the maximum vertex transmission of $G$. Then the difference between the maximum transmission and the generalized distance spectral radius of $G$ is $(1-\alpha)\tau(G)$.
In [3], Atik and Panigrahi introduced a class of graphs called $DVDR$ graphs. Let $G$ be a connected graph with order $n$. If there exist a vertex $v$ in $G$ such that the degree $d(v)=n-1$ and $G-v$ is an regular graph, then we call $G$ to be a distinguished vertex deleted regular graph ($DVDR$). The vertex $v$ with $d(v)=n-1$ is said to be distinguished vertex of graph $G$. In addition, we say $G$ is an $r$-$DVDR$ graph if $G$-$v$ is $r$-$regular$.
Let $G_{0}$ be a graph which attains the minimum of $Tr_{max}(G)-\mu_{\alpha}(G)$ among all connected non-transmission regular graph $G$ of order $n$. Since $G_{0}$ is a connected graph, then it follows from the Perron Frobenius Theorem [10] that $\mu_{\alpha}(G_{0})$ is an eigenvalue of $D_{\alpha}(G_{0})$, and there is a unique positive unit eigenvector corresponding to $\mu_{\alpha}(G_{0})$. Let $X=(x_{1},x_{2},\cdots,x_{n})$ be the unique positive unit eigenvector corresponding to $\mu_{\alpha}(G_{0})$. To simplify the writing process, let $x_{max}=max\{x_{i}|1\leq i\leq n\}$ and $x_{min}=min\{{x_{i}|1\leq i\leq n\}}$. Let $Tr_{max}=max\{Tr(v)|v\in V(G_{0})\}$ and $W=W(G_{0})=\frac{1}{2}\sum\limits_{w\in V(G_{0})}Tr_{w}$. By a simple calculation, we get
$Tr_{max}-\mu_{\alpha}(G_{0})$ $=X^{T}(Tr_{max}I_{n})X-X^{T}D_{\alpha}(G_{0})X$ (1)
$=Tr_{max}-2(1-\alpha)\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})x_{i}x_{j}-\alpha\sum\limits_{i=1}^{n}Tr_{i}x_{i}^{2}$
$\geq Tr_{max}-2(1-\alpha)\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})x_{i}x_{j}-\alpha Tr_{max}\sum\limits_{i=1}^{n}x_{i}^{2}$
$=(1-\alpha)[Tr_{max}-2\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})x_{i}x_{j}]$
$=(1-\alpha)[Tr_{max}\sum\limits_{i=1}^{n}x_{i}^{2}-\sum\limits_{i=1}^{n}Tr_{i}x_{i}^{2}+\sum\limits_{i=1}^{n}Tr_{i}x_{i}^{2}-2\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})x_{i}x_{j}]$
$=(1-\alpha)[\sum\limits_{i=1}^{n}(Tr_{max}-Tr_{i})x_{i}^{2}+\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})(x_{i}-x_{j})^{2}]$
$\geq(1-\alpha)[(nTr_{max}-2W)x_{min}^{2}+\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})(x_{i}-x_{j})^{2}],$
where $I_{n}$ is an identity matrix of order $n$.
@@
Let $n=2r-1$ and $K=K_{1,\underbrace{2,2,\cdots,2}_{r-1}}$ be the complete $r$-$partite$ graph on $n$ vertices which all subsets contain two vertices except one subset contains only one vertex $v$ with $d(v)=n-1$. Then
@@
By Lemma 2.3, we know that $\mu_{\alpha}(K)$ is equal to the largest eigenvalue of $M$. Hence, we have
@@
Let $G$ be an $(n-4)$-$DVDR$ graph of order $n$. Let $v$ be the distinguished vertex with $d(v)=n-1$. Then
@@
By Lemma 2.3, we know that $\mu_{\alpha}(G)$ is equal to the largest eigenvalue of $N$. Hence, we have
@@
It is obvious that if $n$ is odd, then $(1-\alpha)\tau_{n}=(1-\alpha)\tau(K_{1,2,2,\cdots,2})$; if $n$ is even, then $(1-\alpha)\tau_{n}=(1-\alpha)\tau(G)$ for any $(n-4)$-$DVDR$ graph $G$. Noting that the above graphs are both connected non-transmission regular graphs and $G_{0}$ be a graph which attains the minimum of $(1-\alpha)\tau(G)$ among all connected non-transmission regular graph $G$ of order $n$. Hence, we have $\tau(G_{0})\leq\tau(G)$. Specifically, $\tau(G_{0})\leq\tau_{n}$.
Theorem 2.6.
Let $G$ be a connected non-transmission regular graph of order $n$. (1) If $n$ is odd and $\alpha\in[0,1)$, then
equality holds if and only if $G\cong K_{1,2,2,\cdots,2}.$ (2) If $n$ is even and $\alpha\in[0,1)$, then
equality holds if and only if $G$ is isomorphic to any $(n-4)$-$DVDR$ graph.
@@
According to the previous note on $G_{0}$, let $u$ and $v$ be two vertices of $G_{0}$ such that $x_{u}=x_{max}$ and $x_{v}=x_{min}$, respectively. Let $\mu_{\alpha}=\mu_{\alpha}(G_{0})$. From the $u$-th equation of $D_{\alpha}(G_{0})X=\mu_{\alpha}X$, we have
@@
Based on the above result, we have
@@
In addition, according to the properties of $G_{0}$, $\tau(G_{0})=\frac{Tr_{\max}-\mu_{\alpha}}{1-\alpha}\leq\tau_{n}$. Hence, we obtain
@@
Claim 1: $\tau_{n}<1$.
Proof: From (4), we consider two cases:
(1) If $n$ is odd, then $\rho_{n}$=1. Thus, $\tau_{n}=\frac{(1-\alpha)n+1-\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2(1-\alpha)}$. We take $\tau_{n}(1)=\tau_{n}$ under these conditions. If $\tau_{n}(1)\geq 1$, by a simple calculation, we get $(1-\alpha)^{2}(n-2)^{2}\geq(1-\alpha)^{2}n^{2}$, this is a contradiction.
(2) If $n$ is even, then $\rho_{n}$=2. Thus, $\tau_{n}=\frac{(1-\alpha)n+2-\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2(1-\alpha)}$. We take $\tau_{n}(2)=\tau_{n}$ under these conditions. Using the same method as (1), we get $\tau_{n}(2)<1$.
Hence, $\tau_{n}<1$, we get
@@
Claim 2: $nTr_{max}-2W=\rho_{n}$.
Proof: Note that $D_{\alpha}(G_{0})X=\mu_{\alpha}X$, we have $\mu_{\alpha}(1,1,\cdots,1)X=(1,1,\cdots,1)D_{\alpha}(G)X$. Furthermore,
@@
Thus, we have
$(Tr_{max}-\mu_{\alpha})\sum_{w\in V(G_{0})}x_{w}=\sum_{w\in V(G_{0})}(Tr_{max}-Tr_{w})x_{w}\\ \geq(nTr_{max}-2W)x_{min}.$ (9)
Therefore, we get that
$nTr_{max}-2W\leq\frac{(Tr_{max}-\mu_{\alpha})\sum\limits_{w\in V(G_{0})}x_{w}}{x_{min}}.\$ (10)
By (8), since $Tr_{max}-\mu_{\alpha}=(1-\alpha)\tau(G_{0})\leq(1-\alpha)\tau_{n}$ and $\sum\limits_{w\in V(G_{0})}x_{w}\leq x_{min}+(n-1)x_{max}$, we have that
$nTr_{max}-2W$ $\leq\frac{(Tr_{max}-\mu_{\alpha})\sum_{w\in V(G_{0})}x_{w}}{x_{min}}$
@@
Then we will check that $nTr_{max}-2W\geq\rho_{n}$. Note that $nTr_{max}-2W$ is a positive integer. If $n$ is odd, we have $\rho_{n}=1$, it is obvious that $nTr_{max}-2W\geq 1=\rho_{n}$. If $n$ is even, we have $\rho_{n}=2$ and $nTr_{max}-2W$ is an even positive integer. Thus we have $nTr_{max}-2W\geq 2=\rho_{n}$. Thus, we have $nTr_{max}-2W=\rho_{n}$. Claim 3: $Tr_{v}<Tr_{max}$ and $Tr_{u}>Tr_{min}$.
Proof: Since $D_{0}(G)X=\mu_{\alpha}X$, from the $v$-th equation, we get
@@
hence $Tr_{v}\leq\mu_{\alpha}$. Since $\mu_{\alpha}<Tr_{max}$, thus $Tr_{v}<Tr_{max}$. Similarly, we get that $Tr_{u}>Tr_{min}$. Claim 4: $Tr_{max}-\mu_{\alpha}\geq\frac{\rho_{n}}{1+(n-1)\frac{x_{max}}{x_{min}}}.$ Proof: It follows from (9) and Claim 2, we have that
$(Tr_{max}-\mu_{\alpha})\sum_{w\in V(G_{0})}x_{w}=\sum_{w\in V(G_{0})}(Tr_{max}-Tr_{w})x_{w}\\ \geq(nTr_{max}-2W)x_{min}=\rho_{n}x_{min}.$ (12)
Hence, we have
@@
Since $\frac{x_{max}}{x_{min}}\leq\frac{1}{1-\tau_{n}}$, we have

### v2

@@ Abstract
The largest eigenvalue of $D_{\alpha}(G)$ is called the $D_{\alpha}$ spectral radius or generalized distance spectral radius of $G$, denoted by $\mu_{\alpha}(G)$. In this paper, we establish a lower bound on the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius of non-transmission regular graphs, and we also characterize the extremal graphs attaining the bound.
Key Words: Generalized distance matrix, Spectral radius, Non-transmission regular graph, Lower bound.
AMS Subject Classification (2020): 05C50,15A18
@@ 1 Introduction
Let $G=(V(G),E(G))$ be a graph with vertex set $V(G)=\{v_{1},v_{2},$ $\ldots,v_{n}\}$ and edge set $E(G)$. The number of vertices $|V(G)|=n$ is the order of $G$ and the number of edges $|E(G)|=m$ is the size of $G$. For any vertex $v_{i}\in V(G)$, $N(v_{i})=\{v_{j}:(v_{i},v_{j})\in E(G)\}$ is called the set of neighbors of $v_{i}$. The degree $d(v_{i})$ of a vertex $v_{i}$ in $G$ is the number of edges of $G$ incident with $v_{i}$, i.e. , $d(v_{i})=|N(v_{i})|$. Let $K_{n}$ denote the complete graph of order $n$. Throughout this paper all graphs considered are simple and connected.
@@
The distance between two vertices $v_{i}$ and $v_{j}$, denoted by $d(v_{i},v_{j})$ or simply by $d_{ij}$, is the length of a shortest path from $v_{i}$ to $v_{j}$ in $G$. The diameter of $G$, denoted by $diam(G)$, is the maximum distance between any two vertices of $G$. For any vertex $v_{i}\in V(G)$, the transmission of $v_{i}$ in $G$, denoted by $Tr_{i}$, is the sum of distances from $v_{i}$ to all other vertices of $G$, that is, $Tr_{i}=\sum\limits_{k=1}^{n}d_{ik}$. We use $Tr_{max}(G)$ to denote the maximum vertex transmission of $G$ (or simply $Tr_{max}$ if it is clear from the context). For a connected graph $G$, if $Tr_{1}=Tr_{2}=\cdots=Tr_{n}=r$, we say that $G$ is $r$-transmission regular; otherwise, $G$ is non-transmission regular. The sum of all distances between all unordered pairs of vertices in $G$ is called the Wiener index of graph $G$, denoted by $W(G)$, i.e., $W(G)=\frac{1}{2}\sum\limits_{i=1}^{n}Tr_{i}$.
@@
For a connected graph $G$, the distance matrix of $G$ is denoted by $D(G)=(d_{ij})_{n\times n}$. Let $Tr(G)=\textrm{diag}(Tr_{1},Tr_{2},\ldots,Tr_{n})$ be the diagonal matrix with vertex transmissions of $G$. With the development of the matrix theory of graphs, Aouchiche and Hansen [1] defined the distance signless Laplacian matrix of $G$ as $D^{Q}(G)=Tr(G)+D(G)$. The largest distance (distance signless Laplacian) eigenvalue of $G$ is called the distance (distance signless Laplacian) spectral radius of $G$, denoted by $\mu(G)$ ($\mu_{Q}(G)$).
@@
So far, the distance spectral radius and distance signless Laplacian spectral radius of a connected graph have been investigated extensively. For example, in [20], Nath and Paul determined the unique graph with the minimum distance spectral radius among all connected bipartite graphs of order $n$ with a given matching number. In [13], Lin and Lu determined the graphs with the minimum (maximum) distance signless Laplacian spectral radius among all connected graphs with fixed order and clique number. In [15], Lin and Zhou determined the unique trees with the second and third maximum distance signless Laplacian spectral radius, respectively, and they also determined the unique connected graph with the second maximum distance signless Laplacian spectral radius. In [27], Zhang et al. determined the minimum distance spectral radius of graphs among all $n$-vertex graphs with given connectivity and independence number, and characterized the corresponding extremal graph. For more results about distance and distance signless Laplacian spectral radius of graphs, readers may refer to [2, 3, 10, 23, 24, 26] and the references therein.
@@
In [7], Cui et al. proposed to study the convex linear combinations of $Tr(G)$ and $D(G)$, denoted by $D_{\alpha}(G)$ and defined as
@@
We call $D_{\alpha}(G)$ as the generalized distance matrix of $G$. It is easy to see $D_{0}(G)=D(G)$, $D_{1}(G)=Tr(G)$ and $2D_{\frac{1}{2}}(G)=D^{Q}(G)$, thus, the matrix $D_{\alpha}(G)$ extends both $D(G)$ and $D^{Q}(G)$. Therefore, it is very important to study the generalized distance matrix. The largest eigenvalue of $D_{\alpha}(G)$ is called the $D_{\alpha}$ spectral radius or the generalized distance spectral radius of $G$, denoted by $\mu_{\alpha}(G)$.
Recently, the generalized distance spectral radius of graphs has been studied in some papers. In [8], Cui et al. characterized the unique graph with the minimum generalized distance spectral radius among the connected graphs with fixed chromatic number. In [9], Diaz et al. presented some extremal results on the spectral radius of generalized distance matrix that generalize previous results on the spectral radii of the distance matrix and distance signless Laplacian matrix. In [14], Lin et al. gave several graph transformations on the $D_{\alpha}$ spectral radius of graphs. There are also some other interesting results on the $D_{\alpha}$ spectral radius of graphs which can be found in [4, 9, 19, 21].
In the theory of graph spectra, there is a classic problem which is to establish upper or lower bounds for the related spectral radius of a given class of graphs. For a graph $G$, the adjacent spectral radius of $G$ is denoted by $\rho(G)$. It is well-known that the maximum degree $\Delta(G)$ is no less than $\rho(G)$. In addition, $\Delta(G)=\rho(G)$ if and only if $G$ is a regular graph. Thus, $\Delta(G)-\rho(G)$ has been considered as a measure of irregularity for a graph $G$ and many interesting results regarding its estimation have been obtained, see [5, 6, 16, 22, 28].
By analogy with adjacency matrices, it is natural to consider the difference between the maximum vertex transmission and the distance spectral radius of non-transmission regular graphs. Motivated by some known estimates on $\Delta(G)-\rho(G)$ and some methods used in estimating them, in 2018, Liu et al. [18] posed and considered the following questions: How small can $Tr_{max}(G)-\mu(G)$ and $2Tr_{max}(G)-\mu_{Q}(G)$ be when $G$ is non-transmission regular? In [18], Liu et al. posed a conjecture about the difference between the maximum vertex transmission and the distance spectral radius of non-transmission regular graphs. In [17], Liu et al. proved the conjecture proposed by Liu et al. [18] and they gave a lower bound for the difference between the maximum vertex transmission and the distance spectral radius of non-transmission regular graphs. In [12], Lan and Liu gave a sharp lower bound of $2Tr_{max}-\mu_{Q}(G)$ among all $n$-vertex connected graphs, and characterized the extremal graphs.
In this paper we decide to study the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius of connected non-transmission regular graphs, we establish a lower bound on the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius of non-transmission regular graphs, and we also characterize the extremal graphs attaining the bound. Our main result is as follows.
Theorem 1.1.
Let $G$ be a connected non-transmission regular graph of order $n$.
(1) If $n$ is odd and $\alpha\in[0,1)$, then
equality holds if and only if $G\cong K_{1,2,2,\cdots,2}$ (the definition see Lemma 2.4).
(2) If $n$ is even and $\alpha\in[0,1)$, then
equality holds if and only if $G$ is isomorphic to any (n4)(n-4)-DVDRDVDR graph (the definition see Page 4).
2 Preliminaries
In this section, we will introduce some important concepts and useful lemmas which will be used in the following.
@@ 2 Preliminaries
([11]) Let $M$ be an irreducible and nonnegative square matrix of order $n$. Let $\rho(M)$ be the spectral radius of $M$. Then the following results hold.
@@
Equitable quotient matrix has significant applications in calculating the eigenvalues of matrices associated with graphs. Next, we will introduce the definition of the equitable quotient matrix and its some properties.
@@
where the blocks $M_{ij}$ are the $n_{i}\times n_{j}$ matrices for any $1\leq i,j\leq t$ and $n=n_{1}+n_{2}+\cdots+n_{t}.$ For $1\leq i,j\leq t$, let $b_{ij}$ denote the average row sum of $M_{ij}$, i.e. $b_{ij}$ is the sum of all entries in $M_{ij}$ divided by the number of rows. Then $B(M)=(b_{ij})$(or simply $B$) is called the quotient matrix of $M$. If, in addition, for each pair $i$, $j$, $M_{ij}$ has a constant row sum, then $B$ is called the equitable quotient matrix of $M$.
@@
([25]) Let $B$ be the equitable quotient matrix of $M$, where $M$ is defined as in Definition $2.$ In addition, let $M$ be a nonnegative matrix, and $\rho(B)$ and $\rho(M)$ denote the spectral radii of $B$ and $M$, respectively. Then $\rho(B)=\rho(M)$.
For a connected graph $G$ with order $n$, take $\tau(G)=\frac{Tr_{max}(G)-\mu_{\alpha}(G)}{1-\alpha}$, where $Tr_{max}(G)$ is the maximum vertex transmission of $G$. Then
$(1-\alpha)\tau(G)=Tr_{max}(G)-\mu_{\alpha}(G).$
In [3], Atik and Panigrahi introduced a class of graphs called $DVDR$ graph. Let $G$ be a connected graph with order $n$. If there exist a vertex $v$ in $G$ such that the degree of $v$ is n1n-1 and GvG-v is an regular graph, then we call $G$ to be a distinguished vertex deleted regular graph ($DVDR$, see Figure 1). The vertex $v$ with d(v)=n1d(v)=n-1 is said to be a distinguished vertex of a graph $G$. In addition, we say $G$ is an $r$-DVDRDVDR graph if GvG-v is $r$-regular.
Figure 1: Some examples of $DVDR$ graphs
@@
Let n=2r1n=2r-1 and $K=K_{1,\underbrace{\scriptstyle 2,2,\cdots,2}_{r-1}}$ be the complete $r$-partite graph on $n$ vertices which all subsets contain two vertices except one subset contains only one vertex $v$ with d(v)=n1d(v)=n-1. Then
@@
By Lemma 2.3, we know that $\mu_{\alpha}(K)$ is equal to the largest eigenvalue of $M$. Hence
@@
Let $G$ be an (n4)(n-4)-DVDRDVDR graph of order $n$. Let $v$ be the distinguished vertex with d(v)=n1d(v)=n-1. Then
@@
By Lemma 2.3, we know that $\mu_{\alpha}(G)$ is equal to the largest eigenvalue of $N$. Hence
@@
It is obvious that if $n$ is odd, then $(1-\alpha)\tau_{n}=(1-\alpha)\tau(K_{1,2,2,\cdots,2})$; if $n$ is even, then $(1-\alpha)\tau_{n}=(1-\alpha)\tau(G)$ for any (n4)(n-4)-DVDRDVDR graph $G$.
Let $G_{0}$ be a graph which attains the minimum of $Tr_{max}(G)-\mu_{\alpha}(G)$ among all connected non-transmission regular graph $G$ of order $n$. Noting that the graphs in Lemma 2.4 and Lemma 2.5 are both connected non-transmission regular graphs and $G_{0}$ be a graph such that attains the minimum of $(1-\alpha)\tau(G)$ among all connected non-transmission regular graph $G$ of order $n$. Hence $\tau(G_{0})\leq\tau(G)$. Specifically, $\tau(G_{0})\leq\tau_{n}$.
To simplify the proof, we first provide some notations that will be used in the proof. Since $G_{0}$ is a connected graph, then it follows from the Perron Frobenius Theorem [11] that $\mu_{\alpha}(G_{0})$ is an eigenvalue of $D_{\alpha}(G_{0})$, and there is a unique positive unit eigenvector corresponding to $\mu_{\alpha}(G_{0})$. We use $X=(x_{1},x_{2},\cdots,x_{n})^{T}$ to denote the positive unit eigenvector corresponding to $\mu_{\alpha}(G_{0})$. To simplify the writing process, let $x_{max}=\max\{x_{i}|1\leq i\leq n\}$ and $x_{min}=\min\{{x_{i}|1\leq i\leq n\}}$. Let $Tr_{max}=\max\{Tr(v)|v\in V(G_{0})\}$ and $W=W(G_{0})=\frac{1}{2}\sum\limits_{w\in V(G_{0})}Tr_{w}$.
In the following sections, we will provide the proof of Theorem 1.1.
3 Proof of Theorem 1.1
@@
According to the previous note on $G_{0}$, let $u$ and $v$ be two vertices of $G_{0}$ that satisfy $x_{u}=x_{max}$ and $x_{v}=x_{min}$, respectively. For convenience, let $\mu_{\alpha}=\mu_{\alpha}(G_{0})$. From the $u$-th equation of $D_{\alpha}(G_{0})X=\mu_{\alpha}X$, we have
@@
Based on the above result, then
@@
In addition, according to the properties of $G_{0}$, it is shown that
$\tau(G_{0})=\frac{Tr_{\max}-\mu_{\alpha}}{1-\alpha}\leq\tau_{n}.$ (7)
Hence, we obtain
@@
Claim 1. $\tau_{n}<1$.
Proof. From (5) and $\alpha\in[0,1)$, we consider two cases as follow:
Case 1. If $n$ is odd, then $\rho_{n}$=1. Thus, $\tau_{n}=\frac{(1-\alpha)n+1-\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2(1-\alpha)}$. We take $\tau_{n}(1)=\tau_{n}$ under this condition. If $\tau_{n}(1)\geq 1$, by a simple calculation, we get $(1-\alpha)^{2}(n-2)^{2}\geq(1-\alpha)^{2}n^{2}$, this is a contradiction.
Case 2. If $n$ is even, then $\rho_{n}$=2. Thus, $\tau_{n}=\frac{(1-\alpha)n+2-\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2(1-\alpha)}$. We take $\tau_{n}(2)=\tau_{n}$ under this condition. Using the same method as Case 1, we have $\tau_{n}(2)<1$.
Hence, $\tau_{n}<1$ and
@@
Claim 2. $nTr_{max}-2W=\rho_{n}$ and $\frac{x_{max}}{x_{min}}=\frac{1}{1-\tau_{n}}$.
Proof. Note that $D_{\alpha}(G_{0})X=\mu_{\alpha}X$, we have $\mu_{\alpha}(1,1,\cdots,1)X=(1,1,\cdots,1)D_{\alpha}(G_{0})X$. Furthermore,
@@
Thus,
$(Tr_{max}-\mu_{\alpha})\sum_{w\in V(G_{0})}x_{w}=\sum_{w\in V(G_{0})}(Tr_{max}-Tr_{w})x_{w}\\ \geq(nTr_{max}-2W)x_{min},$ (9)
and
$nTr_{max}-2W\leq\frac{(Tr_{max}-\mu_{\alpha})\sum\limits_{w\in V(G_{0})}x_{w}}{x_{min}}.\\$ (10)
Observe that $Tr_{max}-\mu_{\alpha}=(1-\alpha)\tau(G_{0})\leq(1-\alpha)\tau_{n}$ and $\sum\limits_{w\in V(G_{0})}x_{w}\leq x_{min}+(n-1)x_{max}$, combining with (8), we obtain
$nTr_{max}-2W$ $\leq\frac{(Tr_{max}-\mu_{\alpha})\sum_{w\in V(G_{0})}x_{w}}{x_{min}}$ (11)
@@
In what follows, we will check that $nTr_{max}-2W\geq\rho_{n}$. Note that $nTr_{max}-2W$ is a positive integer. If $n$ is odd, we have $\rho_{n}=1$, it is obvious that $nTr_{max}-2W\geq 1=\rho_{n}$. If $n$ is even, we have $\rho_{n}=2$ and $nTr_{max}-2W$ is an even positive integer. Thus $nTr_{max}-2W\geq 2=\rho_{n}$.
Hence, it can be concluded that $nTr_{max}-2W=\rho_{n}$. Moreover, all inequalities in (11) must be equalities. Thus
$\frac{x_{max}}{x_{min}}=\frac{1}{1-\tau_{n}}.$
Claim 3. $Tr_{v}<Tr_{max}$ and $Tr_{u}>Tr_{min}$.
Proof. Since $D_{\alpha}(G_{0})X=\mu_{\alpha}X$, from the $v$-th equation, we get
@@
i.e.,
$Tr_{v}\leq\mu_{\alpha}.$
Combining with $\mu_{\alpha}<Tr_{max}$, then $Tr_{v}<Tr_{max}$.
Similarly, we obtain $Tr_{u}>Tr_{min}$.
Claim 4. $Tr_{max}-\mu_{\alpha}=(1-\alpha)\tau_{n}.$
Proof. It follows from (9) and Claim 2, we have
$(Tr_{max}-\mu_{\alpha})\sum_{w\in V(G_{0})}x_{w}\geq(nTr_{max}-2W)x_{min}=\rho_{n}x_{min}.$ (13)
Hence
@@
From Claim 2, we know that $\frac{x_{max}}{x_{min}}=\frac{1}{1-\tau_{n}}$, then